A vending treasury is purchased for RM 2,800 in cash with RM 700 down payment and fourteen monthly payments. If the interest calculated is 10% per annum, using the fixed rate equation calculation find:


I. Total interest charged


Secondly. monthly payment



1
Expert's answer
2022-02-01T13:37:43-0500

Solution:

Amount need to borrow = 2800-700 = 2100

Monthly Rate "= \dfrac{10\%}{12} = 0.8333\%"

Monthly Payment "= \dfrac{Amount\ need\ to\ borrow}{\dfrac{1-(1+r)^{-n}}r}"

"= \dfrac{2100}{\dfrac{1-(1+0.8333\%)^{-14}}{0.8333\%}}"

=$159.54

Total interest charged = Amount to be paid in total - Amount Borrowed

"=14\times159.54-2100"

=$133.60